OpenStudy (anonymous):

Hi, I am doing a Related Rates problem and I got stuck on finding the derivative with respect to time of the equation for the volume of a cylinder: V = (pi)(r^2)h

OpenStudy (anonymous):

V = (pi)(r^2)h this does not depend on time

OpenStudy (anonymous):

so derivative is =0

OpenStudy (anonymous):

it doesn't? hmm.. here lemme write out the problem because i thought this would be the right equation to use

OpenStudy (anonymous):

piston is seated at the top of a cylindrical chamber with radius 5 cm when ti starts moving into the chamber at a constant speed of 8 cm/s. What is the rate of change of the volume of the cylinder when the piston is 6cm from the base of the chamber?

OpenStudy (anonymous):

so i copied this in my notes: \[V = \pi r^2h\] \[dV/dt = d/dt (\pi r^2h)\] \[= \pi r^2dh/dt\] i don't get how she went from the 2nd step to the 3rd

OpenStudy (anonymous):

h is function of time. It is h(t)=\(h_0\)-8t. Where \(h_0\) is cylinder height. So dh/dt=-8

OpenStudy (anonymous):

and rate of change of volumen will be \(\pi\)25(-8)=?